The cost and production data in 2006 for a manufacturing company in Tennessee is provided in the table below.  The selling price is fixed adding a 40% mark- up on the production cost. The selling price/unit is equal to?The cost and production data in 2006 for a manufacturing company in Tennessee is provided in the table below. Material and Pa
Answer #1

The total product cost = Material and Parts cost + Labor costs + overhead cost

= 1,500,000 + 2,000,000 + 2,500,000

= 6,000,000

Now, the product cost per unit of product (before applying the markup value)  = 6,000,000 / 500,000

= 12

Selling price = 6,000,000 + 40% of 6,000,000

= 8,400,000

Answer : The selling price per unit = (8,400,000 / 500,000) = 16.8